OpenStudy (anonymous):

solve by substitution −2x−y=−9 5x−2y=18

OpenStudy (anonymous):

from (1) we get 5x-2(-2x+9)=18 9x=36 x=4

OpenStudy (anonymous):

y=1

OpenStudy (anonymous):

y = 9-2x substitute to 5x-2y=18 5x-2(9-2x)=18 5x+4x-18=18 9x=36 x=4 substitute to y =9-2x y=9-2(4) =1